# Verification Learning: Make Unsupervised Neuro-Symbolic System Feasible

### v1

@@ Abstract
The current Neuro-Symbolic (NeSy) Learning paradigm suffers from an over-reliance on labeled data. If we completely disregard labels, it leads to less symbol information, a larger solution space, and more shortcuts—issues that current Nesy systems cannot resolve. This paper introduces a novel learning paradigm, Verification Learning (VL), which addresses this challenge by transforming the label-based reasoning process in Nesy into a label-free verification process. VL achieves excellent learning results solely by relying on unlabeled data and a function that verifies whether the current predictions conform to the rules. We formalize this problem as a Constraint Optimization Problem (COP) and propose a Dynamic combinatorial Sorting (DCS) algorithm that accelerates the solution by reducing verification attempts, effectively lowering computational costs to the level of a Constraint Satisfaction Problem (CSP). To further enhance performance, we introduce a prior alignment method to address potential shortcuts. Our theoretical analysis points out which tasks in Nesy systems can be completed without labels and explains why rules can replace infinite labels, such as in addition, for some tasks, while for others, like Sudoku, the rules have no effect. We validate the proposed framework through several fully unsupervised tasks including addition, sort, match, and chess, each showing significant performance and efficiency improvements.
@@ 1 Introduction
Human cognition operates through a dual-system framework: System 1 (intuitive processing) enables rapid, experience-based responses to simple tasks, while System 2 (deliberative processing) engages slow, knowledge-intensive reasoning for complex ones. This cognitive architecture mirrors the dichotomy in artificial intelligence (AI) between data-driven and rule-driven learning paradigms. Rule-based symbolic learning systems have been replaced by data-driven neural learning systems due to the high cost of maintenance and search. However, with the continuous iteration of neural network models, they still perform poorly on reasoning tasks. Consequently, an increasing number of research efforts are now focused on Nesy systems capable of integrating perception and reasoning.
In the Nesy paradigm, a machine learning model $f$ establishes a mapping between inputs $X$ and symbolic representations $S$, i.e., $S=f(X)$. Then, using a knowledge base $KB$ and $S$, to infer the label $Y$, i.e., $KB,S\models Y$. $S$ is unknown and needs to be predicted based on both $X$ and $Y$, ensuring consistency between the learning and reasoning processes. Currently, the real-world application of Nesy remains limited. For simple tasks, good results can be achieved without a symbolic system. For complex tasks, the required data grows exponentially. However, most current Nesy algorithms still require labels $Y$ of the same scale as the input data $X$. The difficulty in bridging the gap between supervised Nesy and unsupervised Nesy remains significant.
@@
Due to the aforementioned issues, applying previous approaches for unsupervised Nesy is largely infeasible. To address these problems, we propose a new paradigm called Verification Learning (VL). In an unsupervised setting, where a knowledge module cannot perform reasoning starting from the label $Y$, we instead replace the reasoning process with a constraint satisfaction verification on $S$ drawn from the solution space. Using satisfiability verification instead of logical reasoning not only resolves the problem of a lack of starting points, but also bypasses challenges inherent in logical reasoning, such as incompleteness, infinite recursion, and high computational complexity. Additionally, this approach is not tied to a specific logical form, allowing more general verification functions to replace the complex logic programming. Moreover, because it does not depend on label $Y$, the VL paradigm supports test time corrections, ensuring that the solutions output by the model conform to the specified constraints.
VL’s learning process corresponds to a Constraint Optimization Problem (COP), where the objective is to find the solution that maximizes the optimization goal among all feasible solutions that satisfy the constraints. If the verification of solutions requires exhaustively traversing the entire solution space, this becomes computationally expensive. However, if we can sort the solutions in the solution space according to their optimization goal values and verify them in order dynamically, we can ensure that the first valid solution we verify is the optimal one. Initially, this sorting process had exponential complexity, but we introduced an algorithm called Dynamic Combinatorial Sorting (DCS). This algorithm maintains a heap structure with low computational overhead, dynamically tracking the solution with the highest optimization goal value among unverified solutions. By doing so, we guarantee that the first valid solution we verify is the optimal one, thus reducing the COP problem’s complexity to a similar level as the CSP. Initially, we implemented DCS under the assumption of independence and later extended it to handle some cases where independence does not hold but monotonicity holds, making it still applicable. This reveals that, in Nesy, monotonicity is a more relaxed yet crucial property compared to independence.
@@ 2.1 Neuro-Symbolic Learning
Research on Nesy can generally be classified into three types(Yu et al., 2023): Reasoning to learning, which focuses on constructing network architectures(Mao et al., 2019) or loss functions based on rules, emphasizing the mapping from $X$ to $S$; Learning to reasoning, where learning methods are employed to identify reasoning paths(Mao et al., 2019) or accelerate the search process(Wang et al., 2018), focusing on the reasoning from $S$ to $Y$; Learning-reasoning: which emphasizes the interaction between the two components and where $S$ is unknown and needs to be obtained based on both $X$ and $Y$, ensuring consistency between the learning and reasoning processes. DeepProblog introduces probabilistic reasoning by incorporating probabilities predicted by neural networks into Prolog programs(Manhaeve et al., 2018), while DeepStochlog, based on DeepProlog, uses Stochastic Definite Clause Grammars for stochastic reasoning(Winters et al., 2022). NeurASP, on the other hand, uses Answer Set Programming as the knowledge base(Yang et al., 2023). Among the above methods, all tend toward probabilistic reasoning. In contrast, abductive learning (ABL), as an important branch, is dedicated to inferring a definite symbol $S$ from $Y$, ensuring that $S$ aligns with the neural network’s predictions(Dai et al., 2019; Huang et al., 2021, 2020). Ground ABL preprocesses a Ground Truth knowledge base based on ABL to avoid using Prolog search(Cai et al., 2021), while weakly supervised ABL (WSABL) computes loss based on confidence-weighted calculations for different candidate symbols $S$(He et al., 2024).
This concept of label leakage was first introduced in the context of SATNet(Wang et al., 2019) solving visual Sudoku problems(Chang et al., 2020), where the model could directly determine symbolic labels $S$ without relying on the input $X$, instead using only the labels $Y$. After eliminating this leakage, SATNet’s performance on digit recognition completely failed, even dropping to 0%. We’ve found that this phenomenon is widespread in Nesy tasks. This reveals a situation in many tasks where the answer can be obtained without learning, significantly lowering the task’s difficulty and failing to accurately reflect the system’s ability to integrate the two components.
For the optimization problem of Nesy solving, it often involves search problems with exponential complexity. Currently, there are some works that optimize this, such as using gradient-free optimization algorithms(Yu et al., 2016), predicting error locations(Morishita et al., 2023), and designing reward-based search methods using reinforcement learning algorithms(Hu et al., 2025; Li et al., 2020). However, these optimization algorithms cannot guarantee that the solution found is globally optimal.
@@
Recently, (van Krieken et al., 2024) explores the issue of independence in Nesy, pointing out that current Nesy frameworks rely on the assumption of independence. However, in reality, symbols do not satisfy independence always.
@@ 3 From Supervised Reasoning to Unsupervised Verification
Here, $score(S^{\prime})$ represents the probability or weight associated with $S^{\prime}$, and the definition of the score varies across different approaches. In DeepProblog and DeepStochlog, the score is the product of the rule probabilities along the reasoning path. In ABL, the score corresponds to the consistency distance between $f(X)$ and $S^{\prime}$, while in WSABL, it reflects the confidence in $g(X)$.
@@ 4 Solve COP by Dynamic Combinatorial Sorting
Given that $S=[s_{1},\dots,s_{l}]$ consists of $l$ symbols, and considering the global feasible set as a combination of multiple variables, this problem is known as the combinatorial sorting problem. Since the size of the feasible set is $k^{l}$, using classical sorting algorithms would be computationally infeasible. However, when the score satisfies certain properties, we can maintain the score rankings of the next assignment with very low complexity after each assignment. This significantly reduces the computational complexity while ensuring that the verification order remains strictly ordered, guaranteeing the global optimality of the COP solution.
First, we discuss the scenario where the score satisfies independence, and present a combinatorial sorting solution(van Krieken et al., 2024). We then extend the solution to cases where independence is not satisfied but monotonicity is, which provides more relaxed conditions. This highlights that monotonicity is a more crucial property for Nesy tasks, and problems that satisfy monotonicity are just as easy to solve as those that satisfy independence.
@@ 4.1 Independent Case
Under the property of independence, while the greedy rule does not hold (i.e., the $i+1$-th assignment $S_{i+1}$ is not necessarily the successor of the $i$-th assignment $S_{i}$), it is always true that $S_{i+1}$ is a successor of one of the preceding $i$ assignments. Formally, for any $i\in k^{l}$, there exists $1\leq j\leq i$ such that $S_{i+1}\in S_{j}$.
@@
    For $i>1$, we select the highest-scoring successor from the heap, i.e., the one with the maximum successor score $S_{j}$, and then select the highest-scoring assignment from $Suc(S_{j})$ to form $S_{i}$. The successor assignments of $S_{j}$ are updated to $Suc(S_{j})\setminus\{S_{i}\}$, and if non-empty, they are placed back in the heap.
By following this process, we ensure that the assignments are verified in strict score order, guaranteeing that the first solution to pass the verification is the optimal solution for the COP problem. If the first valid solution is ranked $K$, the complexity of maintaining the heap structure is $O(K\log K)$, and the total complexity becomes $O(K(\log K+l\log l+k\log c))$.
@@ 4.2 Non-Independent Case
When the Score satisfies monotonicity, i.e., for any $0<p\leq l$ and any $0<q<k$, there does not exist a $S$ such that modifying $s_{p}$ from $s_{p,q}$ to $s_{p,q+1}$ results in $s_{p^{\prime}}$, and $Suc(S^{\prime})_{p}=(s_{1},\dots,s_{p-1},s_{p^{\prime}},s_{p+1},\dots,s_{l})\in Suc(S^{\prime})$, satisfying $\text{Score}(Suc(S^{\prime})_{p})>\text{Score}(S^{\prime})$. Then for any $i>1$, there exists $0<j<i$, $0<p\leq l$ results in $Suc(S_{j})_{p}\in Suc(S_{j})$. Furthermore, there does not exist any $S^{\prime}\in\mathcal{S}$ such that $S^{\prime}\notin\{S_{1},\dots,S_{i-1}\}$ and $\text{Score}(S^{\prime})>\text{Score}(Suc(S_{j})_{p})$. Therefore, $S_{i}=Suc(S_{j})_{p}$.
@@
If even monotonicity is violated, solving for the optimal $Score(S)$ without traversing all solutions becomes difficult. In such cases, strategies like reinforcement learning or other sampling-based methods can be used to estimate $Score(S)$. However, these methods no longer guarantee a strict order, and thus, cannot ensure the optimal solution to the COP problem.
@@ 5 Mitigate Shortcut Problem by Distribution Alignment
To address this, we propose a distribution alignment strategy. By aligning the model’s output symbol distribution with the natural distribution $P$ of the symbols, we can significantly mitigate this issue. In cases where the natural distribution is unknown, we can use a uniform distribution as a prior. During the initial phase, this ensures that the output distribution is spread out. In later stages, even after removing the influence of the prior distribution, the model’s output distribution will naturally spread out and eventually concentrate on the distribution consistent with the training data. The specific adjustment process is to directly modify the probabilities output by the model to:
@@ 6 Theoretical Study on When Knowledge Can Assist Unsupervised Learning
In fact, the theoretical analysis of unsupervised learning can be divided into two parts. The first part is whether the system has the ability to group samples belonging to the same category together. The second part is whether the system can establish a correspondence between the categories identified by the learner and the labels of the symbols. Together, these two factors determine the upper bound of the generalization error in unsupervised learning. VL effectively addresses both of these issues, with the two modules working together to promote each other’s success.
VL can make the predicted results align with the rule base. So our primary goal is to explore the gap between maintaining consistency with the rule base and being able to distinguish between symbols. This gap is determined by the symmetry of the verification function $V_{KB}$ on the symbol set.
@@
We perform an orbit decomposition of the symmetry group $G$ for the verification function $V_{KB}$. For each symbol $s_{i}$, its orbit is defined as $O_{s_{i}}=\{\sigma(s_{i})|\sigma\in G\}$, which represents all symbols that can be reached by applying the symmetries of $G$ to $s_{i}$. This orbit corresponds to the equivalence symbol of $s_{i}$.
@@
This error, caused by the task itself, cannot be compensated by data, making it a fundamental limitation that cannot be addressed through knowledge validation learning. To better understand these concepts, consider the following examples:
@@
    Sudoku Task: In the Sudoku task, any permutation of numbers within a row, column, or 3x3 grid still satisfies the constraints. This means that all permutations of symbols in Sudoku do not alter the knowledge base. Consequently, the upper bound on task-induced error for Sudoku is $R^{up}_{sudoku}=100\%$, meaning that regardless of data size or model performance, the fully unsupervised Sudoku task could still have a 0% accuracy.
@@
    Chess Task: It is also important to consider cases where a one-way inclusion (not a bijection) occurs. For example, in a chessboard scenario, a Queen can move along diagonal and straight lines, a Rook along straight lines, and a Bishop along diagonals. This leads to situations where a Rook or Bishop could be mapped as a Queen. In such cases, the task becomes unsolvable because it lacks the necessary bijections. However, the presence of a natural distribution $P$ helps correct this issue by ensuring that, the model can distinguish Rooks and Bishops from Queens. So $R^{up}_{chess}=0\%$. Thus, the distribution $P$ extends the domain of solvable problems, enabling the model to handle more complex tasks that would otherwise be unsolvable in an unsupervised setting.
Once the error induced by the task is determined, combining it with the error introduced by the learning process allows us to obtain the upper bound on the generalization error. In unsupervised learning, since it is not possible to distinguish between the symbols in the same orbit, the empirical error should be minimized by selecting the permutation in the symmetry group $G$ that minimizes the error. We define $\hat{R}(f)$ as the current minimal symmetric permutation empirical error.

### v2

@@ Abstract
The current Neuro-Symbolic (NeSy) Learning paradigm suffers from an over-reliance on labeled data, so if we completely disregard labels, it leads to less symbol information, a larger solution space, and more shortcuts—issues that current Nesy systems cannot resolve. This paper introduces a novel learning paradigm, Verification Learning (VL), which addresses this challenge by transforming the label-based reasoning process in Nesy into a label-free verification process. VL achieves excellent learning results solely by relying on unlabeled data and a function that verifies whether the current predictions conform to the rules. We formalize this problem as a Constraint Optimization Problem (COP) and propose a Dynamic Combinatorial Sorting (DCS) algorithm that accelerates the solution by reducing verification attempts, effectively lowering computational costs and introduce a prior alignment method to address potential shortcuts. Our theoretical analysis points out which tasks in Nesy systems can be completed without labels and explains why rules can replace infinite labels for some tasks, while for others the rules have no effect. We validate the proposed framework through several fully unsupervised tasks including addition, sort, match, and chess, each showing significant performance and efficiency improvements.
@@ 1 Introduction
Human cognition operates through a dual-system framework: System 1 (intuitive processing) enables rapid, experience-based responses to simple tasks, while System 2 (deliberative processing) engages slow, knowledge-intensive reasoning for complex ones. This cognitive architecture mirrors the dichotomy in artificial intelligence (AI) between data-driven and rule-driven learning paradigms. In the past, rule-driven symbolic learning systems were less favored compared to data-driven neural learning systems due to their high maintenance and search costs. However, with the continuous iteration of neural network models, they still perform poorly on reasoning tasks. Consequently, an increasing number of research efforts are now focused on Nesy systems capable of integrating data-driven and rule-driven methods.
In the Nesy paradigm, a machine learning model $f$ establishes a mapping between inputs $X$ and symbolic representations $S$, i.e., $S=f(X)$. Then, using a knowledge base $KB$ and $S$, to infer the label $Y$, i.e., $KB,S\models Y$. $S$ is unknown and needs to be predicted based on both $X$ and $Y$, ensuring consistency between the learning and reasoning processes. While the goal is to leverage knowledge to reduce reliance on labeled data, the real-world application of Nesy remains limited. For simple tasks, good results can be achieved without a symbolic system. For complex tasks, the required data grows exponentially but most current Nesy algorithms still require labels $Y$ of the same scale as the input data $X$ which are expensive to acquire. Therefore, there is an urgent need to develop unsupervised Nesy systems.
@@
Figure 1: This example illustrates the differences between unsupervised and supervised neuro-symbolic systems in the addition task. It highlights that unsupervised neuro-symbolic systems exhibit broader applicability but also pose greater learning challenges.
Due to the aforementioned issues, applying previous approaches for unsupervised Nesy is largely infeasible. To address these problems, we propose a new paradigm called Verification Learning (VL). In the VL paradigm, the model’s entire learning process (illustrated in Figure 2) can be completed using only unlabeled data and a knowledge-based verification function. In the unsupervised setting, where a knowledge module cannot perform reasoning starting from the label $Y$, we instead replace the reasoning process with a constraint satisfaction verification on $S$ drawn from the solution space. Using satisfiability verification instead of logical reasoning not only resolves the problem of a lack of starting points, but also bypasses challenges inherent in logical reasoning, such as incompleteness, infinite recursion, and high computational complexity. Additionally, this approach is not tied to a specific logical form, allowing more general verification functions to replace the complex logic programming. Moreover, because it does not depend on label $Y$, the VL paradigm supports test time corrections, ensuring that the solutions output by the model conform to the specified constraints.
Figure 2: This example illustrates how verification learning can complete the entire model training process using only unlabeled data and a knowledge-based verification function.
In the problem-solving process, our goal is to identify the symbolic labels that maximize the alignment between the knowledge base and the machine learning predictions. VL’s learning process corresponds to a Constraint Optimization Problem (COP), where the objective is to find the solution that maximizes the optimization score among all feasible solutions that satisfy the constraints. If the verification of solutions requires exhaustively traversing the entire solution space, this becomes computationally expensive. However, if we can sort the solutions in the solution space according to their scores and verify them in order dynamically, we can ensure that the first valid solution we verify is the optimal one. Initially, this sorting process had exponential complexity, but we introduced an algorithm called Dynamic Combinatorial Sorting (DCS), which is an extension of  (Jia et al., 2025). This algorithm maintains a heap structure with low computational overhead, dynamically tracking the solution with the highest optimization goal value among unverified solutions. By doing so, we guarantee that the first valid solution we verify is the optimal one, thus reducing the COP problem’s complexity to a similar level as the CSP. Initially, we implemented DCS under the assumption of independence and later extended it to handle some cases where independence does not hold but monotonicity holds, making it still applicable. This reveals that, in Nesy, monotonicity is a more relaxed yet crucial property compared to independence.
@@ 2.1 Neuro-Symbolic Learning
Research on Nesy can generally be classified into three types (Yu et al., 2023): Reasoning to learning, which focuses on constructing network architectures (Mao et al., 2019) or loss functions based on rules, emphasizing the mapping from $X$ to $S$; Learning to reasoning, where learning methods are employed to identify reasoning paths (Mao et al., 2019) or accelerate the search process (Wang et al., 2018), focusing on the reasoning from $S$ to $Y$; Learning-reasoning: which emphasizes the interaction between the two components and where $S$ is unknown and needs to be obtained based on both $X$ and $Y$, ensuring consistency between the learning and reasoning processes (Xu et al., 2018; Stammer et al., 2023; Petersen et al., 2021). DeepProblog introduces probabilistic reasoning by incorporating probabilities predicted by neural networks into Prolog programs (Manhaeve et al., 2018), while DeepStochlog, based on DeepProlog, uses Stochastic Definite Clause Grammars for stochastic reasoning (Winters et al., 2022). NeurASP, on the other hand, uses Answer Set Programming as the knowledge base (Yang et al., 2023). Among the above methods, all tend toward probabilistic reasoning. In contrast, abductive learning (ABL), as an important branch, is dedicated to inferring a definite symbol $S$ from $Y$, ensuring that $S$ aligns with the neural network’s predictions (Dai et al., 2019; Huang et al., 2021, 2020; He et al., 2024; Shao et al., 2025; Jia et al., 2025; Hu et al., 2025b). Ground ABL preprocesses a Ground Truth knowledge base based on ABL to avoid using Prolog search (Cai et al., 2021), while A3BL computes loss based on confidence-weighted calculations for different candidate symbols $S$ (He et al., 2024).
The concept of label leakage was first introduced in the context of SATNet (Wang et al., 2019) solving visual Sudoku problems (Chang et al., 2020), where the model could directly determine symbolic labels $S$ without relying on the input $X$, instead using only the labels $Y$. After eliminating this leakage, SATNet’s performance on digit recognition completely failed, even dropping to 0%. We’ve found that this phenomenon is widespread in Nesy tasks. This reveals a situation in many tasks where the answer can be obtained without learning, significantly lowering the task’s difficulty and failing to accurately reflect the system’s ability to integrate the two components.
For the optimization problem of Nesy solving, it often involves search problems with exponential complexity. Currently, there are some works that optimize this, such as using gradient-free optimization algorithms (Yu et al., 2016), predicting error locations (Morishita et al., 2023), and designing reward-based search methods using reinforcement learning algorithms (Hu et al., 2025a; Li et al., 2020). However, these optimization algorithms cannot guarantee that the solution found is globally optimal.
@@
Recently, van Krieken et al. (2024) explores the issue of independence in Nesy, pointing out that current Nesy frameworks rely on the assumption of independence. However, in reality, symbols do not satisfy independence always.
@@ 3 From Supervised Reasoning to Unsupervised Verification
Here, $score(S^{\prime})$ represents the probability or weight associated with $S^{\prime}$, and the definition of the score varies across different approaches. In DeepProblog and DeepStochlog, the score is the product of the rule probabilities along the reasoning path. In ABL, the score corresponds to the consistency distance between $f(X)$ and $S^{\prime}$, while in A3BL, it reflects the confidence in $g(X)$.
@@ 4 Solve COP by Dynamic Combinatorial Sorting
Given that $S=[s_{1},\dots,s_{l}]$ consists of $l$ symbols, and considering the global feasible set as a combination of multiple variables, this problem is known as the combinatorial sorting problem. Since the size of the feasible set is $k^{l}$, using classical sorting algorithms would be computationally infeasible. However, when the score satisfies certain properties, We can infer the $(i+1)_{th}$ ranked solution based on the top $i$ known solutions. In this way, we only need to know the highest-scoring solution to sequentially derive all other solutions, without having to enumerate and sort all of them. This significantly reduces the computational complexity while ensuring that the verification order remains strict, guaranteeing the global optimality of the COP solution.
First, we discuss the scenario where the score satisfies independence (van Krieken et al., 2024), and present a combinatorial sorting solution. We then extend the solution to cases where independence is not satisfied but monotonicity is, which provides more relaxed conditions. This highlights that monotonicity is a more crucial property for Nesy tasks, and problems that satisfy monotonicity are just as easy to solve as those that satisfy independence.
@@ 4.1 Independent Case
Under the property of independence, while the greedy rule does not hold (i.e., the $(i+1)_{th}$ assignment $S_{i+1}$ is not necessarily the successor of the $i_{th}$ assignment $S_{i}$), it is always true that $S_{i+1}$ is a successor of one of the preceding $i$ assignments. Formally, for any $i\in k^{l}$, there exists $1\leq j\leq i$ such that $S_{i+1}\in Suc(S_{j})$.
@@
    For $i>1$, we select the highest-scoring successor from the heap, i.e., the one with the maximum successor score $S_{j}$, and then select the highest-scoring assignment from $Suc(S_{j})$ to form $S_{i}$. We then derive the successors $Suc(S_{i})$, which are then placed in the heap. The successor assignments of $S_{j}$ are updated to $Suc(S_{j})\setminus\{S_{i}\}$, and if non-empty, they are placed back in the heap.
By following this process, we ensure that the assignments are verified in strict score order, guaranteeing that the first solution to pass the verification is the optimal solution for the COP problem. If the first valid solution is ranked $K$, the complexity of maintaining the heap structure is $O(K\log K)$, and the total complexity becomes $O(K(\log K+l\log l+k\log k))$.
@@ 4.2 Non-Independent Case
When the Score satisfies monotonicity, i.e., for any $0<p\leq l$ and any $0<q<k$, there does not exist a $S^{\prime}$ such that modifying $s_{p}$ from $s_{p,q}$ to $s_{p,q+1}$ results in $s_{p^{\prime}}$, and $Suc(S^{\prime})_{p}=(s_{1},\dots,s_{p-1},s_{p^{\prime}},s_{p+1},\dots,s_{l})\in Suc(S^{\prime})$, satisfying $\text{Score}(Suc(S^{\prime})_{p})>\text{Score}(S^{\prime})$, then for any $i>1$, there exists $0<j<i$, $0<p\leq l$ results in $Suc(S_{j})_{p}=(s_{1},\dots,s_{p-1},s_{p^{\prime}},s_{p+1},\dots,s_{l})\in Suc(S_{j})$. Furthermore, there does not exist any $S^{\prime}\in\mathcal{S}$ such that $S^{\prime}\notin\{S_{1},\dots,S_{i-1}\}$ and $\text{Score}(S^{\prime})>\text{Score}(Suc(S_{j})_{p})$. Therefore, $S_{i}=Suc(S_{j})_{p}$.
@@
It is worth noting that when monotonicity is not satisfied, even obtaining the top-ranked solution requires exhaustively searching the entire solution space with exponential complexity. Therefore, the following theorem holds.
Theorem 4.4.
The monotonicity of the score function is a sufficient and necessary condition for the COP problem to admit a general algorithm with sub-exponential complexity.
If even monotonicity is violated, solving for the optimal $Score(S)$ without traversing all solutions becomes impossible. In such cases, strategies like reinforcement learning or other sampling-based methods can be used to estimate $Score(S)$. However, these methods no longer guarantee a strict order, and thus, cannot ensure the optimal solution to the COP problem.
@@ 5 Mitigate Shortcut Problem by Distribution Alignment
To address this, we propose a distribution alignment strategy. By aligning the model’s output symbol distribution with the natural distribution $P$ of the symbols, we can significantly mitigate this issue. In cases where the natural distribution is unknown, we can use a uniform distribution as a prior. During the initial phase, this ensures that the output distribution is spread out. In later stages, even after removing the influence of the prior distribution, the model’s output distribution will eventually concentrate on the distribution consistent with the training data. The specific adjustment process is to directly modify the probabilities output by the model to:
@@ 6 Theoretical Study on When Knowledge Can Assist Unsupervised Learning
In fact, the theoretical analysis of unsupervised learning can be divided into two parts. The first component is the knowledge-induced error, which reflects the system’s ability to establish a correspondence between the categories identified by the learner and the labels of the symbols. The second part is the data-induced error, which reflects the system’s ability to group samples belonging to the same category together. Together, these two factors determine the upper bound of the generalization error in unsupervised learning. VL effectively addresses both of these issues, with the two modules working together to promote each other’s success.
6.1 Knowledge-Induced Error in VL
VL can make the predicted results align with the rule base. So our primary goal is to explore the gap between maintaining consistency with the rule base and being able to distinguish between symbols. This gap is determined by the symmetry of the function $V_{KB}$ on the symbol set.
@@ 6.1 Knowledge-Induced Error in VL
We perform an orbit decomposition of the symmetry group $G$ for the function $V_{KB}$. For each symbol $s_{i}$, its orbit is defined as $O_{s_{i}}=\{\sigma(s_{i})|\sigma\in G\}$, which represents all the equivalence symbols that can be reached by applying the symmetries of $G$ to $s_{i}$.
@@
This error, caused by the task itself, cannot be compensated by data, making it a fundamental limitation that cannot be addressed through knowledge verification. To better understand these concepts, consider the following examples:
@@
    Sudoku Task: In the Sudoku task, any permutation of numbers still satisfies the constraints. This means that all permutations of symbols in Sudoku do not alter the knowledge base. Consequently, the upper bound on task-induced error for Sudoku is $R^{up}_{sudoku}=100\%$, meaning that regardless of data size or model performance, the fully unsupervised Sudoku task could still have a 0% accuracy.
@@
    Chess Task: It is also important to consider cases where a one-way inclusion (not a bijection) occurs. For example, in a chessboard scenario, a Queen can move along diagonal and straight lines, a Rook along straight lines, and a Bishop along diagonals. This leads to situations where a Rook or Bishop could be mapped as a Queen. In such cases, the task becomes unsolvable because it lacks the necessary bijections. However, the presence of a natural distribution $P$ helps correct this issue by ensuring that, the model can distinguish Rooks and Bishops from Queens. So $R^{up}_{chess}=0\%$. Thus, the distribution $P$ extends the domain of solvable problems, enabling the model to handle more complex tasks that would otherwise be unsolvable.
6.2 Data-Induced Error in VL
Once the error induced by the task is determined, combining it with the error introduced by the learning process allows us to obtain the upper bound on the generalization error. In unsupervised learning, since it is not possible to distinguish between the symbols in the same orbit, the empirical error should be minimized by selecting the permutation in the symmetry group $G$. We define $\hat{R}(f)$ as the current minimal symmetric permutation empirical error.